OpenStudy (anonymous):

what is the solution of the equation -3(y-1)+2=-5(y+1)-7?

OpenStudy (anonymous):

-3(y-1)+2=-5(y+1)-7 distribute the -3 and the -5 -3y + 3 + 2 = -5y - 5 - 7 move all the y terms to one side -3y + 5y = -5 - 7 + 2 + 3 combine like terms 2y = -7 solve for y